suppose that you walk at a constant of 6 feet per second. how long do you have to walk to cover a distance of 4 miles ?
round to the nearest tenth.
________minutes

Answer by rfer(16322)

5280/6=880 sec per mi
880*4=3520 sec
3520/60=59 minutes